OpenStudy (anonymous):

Solve the system of equations. 8x – 3y = 1 –2x + 3y = 11 A. (–1, –3) B. (–1, 3) C. (2, 5) D. (5, 2)

OpenStudy (paki):

8x-3y+1=0 -2x+3y-11=0 now solve for x and y...

OpenStudy (anonymous):

i don't know how can u please help please

OpenStudy (anonymous):

the answer is c

OpenStudy (paki):

see... 8x-3y=-1 .... equation 1 -2x+3y=11 .... equation 2 add equation 1 and 2... you will get... x= 10/6 = 5/3... now put this value of "x" in equation 1 or 2... you will get "y" value... hp[e it works now.... @chas242000